find domain of the function
gx= _2_
5-7x

Answer by stanbon(75887)

find domain of the function
gx= _2_
5-7x
=================
That denominator cannot be zero.
It will be zero
if 5-7x = 0
if 7x = 5
if x = 5/7
=========================
Domain: All Real Numbers except x = 5/7
===========================================
Cheers,
Stan H.